Solve the given system of DE by Elimination method. 3) (D - 2)x - y =- et - 3x + (D - 4)y=- 7 et
Answer #1

(0-2)* -Y=-et -34 + (0-4) Y = - 7et le CD-2) 0 – 4) a - (0 - 4)2 =-Get-48+) = 3e -3* + (0-4) y = - Jet - 09 Adding tas we get- La Casa de les Therefore the general solution for a is o n a = Aet+Best+tet And ya CD ~ 2) a te n = D (nett Berntet) = 2(Ae